OpenStudy (anonymous):

i have (sin²x) ÷ (sinx + sinxcosx). I want to get rid of the power. If i do, what would the denominator be?

OpenStudy (anonymous):

how did you intend to get rid of the power?

OpenStudy (anonymous):

\[\frac{ \sin x }{ \sin x + \sin x \cos x }\] you could factor a sin x from the bottom, and that would help

jimthompson5910 (jim_thompson5910):

sin²x is the same as (sinx)*(sinx) sinx + sinxcosx is the same as sinx(1+cosx)

jimthompson5910 (jim_thompson5910):

I'm sure you can see what cancels

OpenStudy (anonymous):

sin^2(x) in the numerator ^

OpenStudy (anonymous):

ohhhok... ill try that and see if i can do it!!